# BANG: Billion-Scale Approximate Nearest Neighbour Search using a Single GPU

### v2

@@ IV-E Prefetching Candidate Nodes
As described in Algorithm 2, candidate node $u^{*}_{i}$ of each query $q_{i}$ will be communicated to the CPU (Line 2) only after the GPU has updated the worklist $\mathcal{L}_{i}$. After that, while the CPU fetches the neighbours (Line 2) and communicates them to the GPU (Line 2), the GPU remains idle, waiting to receive this information. To avoid this delay in starting the current iteration, GPU must be supplied with the neighbour IDs as soon as needed. To address this, instead of sending the candidate nodes at the end of the previous iteration to CPU, we perform an optimization to eagerly identify the next candidate nodes immediately after the neighbour distance calculation (i.e., before even sorting the neighbour list by distance and updating worklist $\mathcal{L}_{i}$).
Our optimization calculates the candidate node $u^{*}_{i}$ eagerly by selecting the nearest neighbour in the new neighbour list $N^{\prime}_{i}$ and the first unvisited node in worklist $\mathcal{L}_{i}$ (sorted by distance to query), then choosing the best between the two. The eager selection occurs just before Line 2 in Algorithm 2. Upon dispatching the eagerly selected candidates to the CPU, the GPU continues executing the remaining tasks of the iteration (sorting the new neighbour list and updating the worklist). Concurrently, the CPU fetches neighbour IDs. That is, Line 2 executes on CPU concurrently with tasks at Line 2 and 2 on GPU. By the completion of the GPU’s remaining tasks, the CPU has already communicated the neighbour IDs. This strategy eliminates GPU idle time, resulting in a throughput improvement (experimentally found to be $\sim 10\%$).
@@ IV-F Parallel Merge Sort
Work-Span Analysis. The work of merging two lists of size $t$ using the parallel merge routine is $O(t\cdot log(t))$ (Section IV-G). The total work $T(n)$ for an array of $n$ numbers follows the recurrence: $T(n)=2\cdot T(n/2)+n\cdot log(n)$, which gives $T(n)=O(n\cdot log^{2}(n))$. Thus, the work of the algorithm is $O(R\cdot log^{2}(R)\cdot\rho)$. Owing to the parallelization scheme, the span of the algorithm is $O(log^{2}(R))$.
@@ IV-G Parallel Merge
The merge routine we detail in this section is employed in the merge sort, as explained in the preceding section. Additionally, it is utilized to merge the sorted neighbour list $\mathcal{N}^{\prime}_{i}$ with the worklist $\mathcal{L}_{i}$ at Line 2 in Algorithm 2.
Figure 3 shows our parallel list merge algorithm on two lists of size four each. Given two sorted lists $L_{in1}$ and $L_{in2}$ with $m$ and $n$ items, where an item $e$ in $L_{in1}$ is at position $p_{1}<m$, and if inserted in $L_{in2}$ would be at position $p_{2}<n$, the position $pos_{Lout}$ of the element in the merged list $L_{out}$ is determined by $pos_{Lout}\leftarrow p_{1}+p_{2}$. We assign one thread to each element. From the thread ID assigned to the item, we can compute the position of the item in its original list. We use binary search to determine where the item should appear in the other list. For example, in Figure 3, the item 28 in $L_{in1}$ is at index 3, determined by the thread ID. In $L_{in2}$, its index is 4, found through binary search. Therefore, the index of 28 in the merged list is 7 (3 + 4). Each thread, assigned to an item, concurrently conducts a binary search in the other list. We keep the lists in the GPU’s shared memory to minimize search latency. Thus, computing the position of each element in the merged list takes $O(\log(m)+\log(n))$ time. Lastly, elements are scattered to their new positions in the merged list, and in the example, 28 is placed at index 7 in the merged list $L_{out}$. This step is also performed in parallel, as each thread writes its item to its unique position in the merged list.
Work-Span Analysis. The work of the algorithm to merge two lists each of size $l$ is $O(l\cdot log(l))$. Owing to the parallelization scheme, the span of the algorithm is $O(log(l))$.
@@ IV-H Re-ranking
We implement the re-ranking step as a separate kernel with optimized thread-block sizes. Re-ranking involves computing the exact (i.e. Euclidean) distance for each query vector with its respective candidate nodes, followed by sorting these candidates by their distance to the query vector and reporting the top-$k$ candidates as nearest neighbours. For a query, its distance from each candidate node is computed in parallel. To sort the candidates according to their exact distance, we use the parallel merge procedure described in Section IV-F. Thanks to the asynchronous data transfers (Section IV-B), the candidates’ base dataset vectors are already available on GPU when this kernel begins. From our experiments, we observed that re-ranking improved the recall by 10-15% for the datasets under consideration.
@@
$O((d\cdot|\mathcal{C}|+|\mathcal{C}|\cdot log^{2}(|\mathcal{C}|))\cdot\rho)$, where $|\mathcal{C}|$ is the maximum number of candidate nodes for a query and $d$ is vector dimension. Owing to the parallelization scheme, the span of re-ranking is $O(d+log^{2}(|\mathcal{C}|))$.
@@ V Implementation Variants
We refer to the implementation described so far in Sections III and IV as BANG. In addition, we provide versions of the implementation specifically optimized for small and medium-sized datasets (having up to 100M points) that can fit in the GPU memory. We discuss these versions below.
@@ V-A In-memory Version
Since the CPU-GPU link stays constantly busy when we store the graph on the CPU, as in BANG, careful task management on the GPU is needed to avoid idle time due to CPU-GPU interdependency. However, following the same approach is unnecessary for smaller graphs that can fit on a GPU. Keeping the entire graph on the GPU, rather than the CPU, eliminates the memory-intensive process of the CPU fetching neighbours from its memory for GPU-supplied candidates, resulting in enhanced throughput. We refer to this version of the implementation as BANG In-memory. The schematic block diagram of this version is shown in Figure 4. There are two main differences from Figure 1: (1) in step 1, we transfer the entire graph from CPU to GPU before the search begins (2) steps 2 and 3 fetch neighbours on the GPU locally through accesses to global memory instead of relying on the CPU.
@@ V-B Exact-distance Version
This variant further optimizes the BANG In-memory version. Instead of using PQ distances within the greedy search loop and subsequently compensating the inaccuracies (resulting due to the use of PQ distances) by the re-ranking step, we directly use exact (Euclidean) distance calculations instead of PQ distances. We refer to this version as BANG Exact-distance. The main differences from Figure 4 are: (1) absence of PQ Distance Table computation, (2) exact distance computation (instead of PQ distance summation), (3) absence of re-ranking step.
@@ VI-A Machine Configuration
Hardware. We conduct our experiments on a machine with 2x Intel Xeon Gold 6326 CPU (with 32 cores), clock-speed 2.90GHz, and 660 GiB DDR4 memory. It houses a 1.41 GHz NVIDIA Ampere A100 GPU with 80 GB global memory and a peak bandwidth of 2039 GB/s. The GPU is connected to the host via a PCIe Gen 4.0 bus having a peak transfer bandwidth of 32 GB/s. We use a single GPU for all the experiments.
Software. The machine runs Ubuntu 22.04.01 (64-bit). We use g++ version 11.3 with -O3, -std=c++11 and -fopenmp flags to compile the C++ code. We use OpenMP for parallelizing the C++ code. The GPU CUDA code is compiled using nvcc version 11.8 with the -O3 flag. BANG source codes are publicly available at [12].
@@ VI-B Datasets
TABLE II: Real-world datasets in our test-suite. (CV Size = Compressed Vectors size)
No. of Data Vector Graph CV
Dataset Type points $d$ Size(GB) Size(GB) Size(GB)
DEEP1B [46] float 1000,000,000 96 384.0 260.0 74.0
SIFT1B [29] uint8 1000,000,000 128 128.0 260.0 74.0
SPACEV1B [13] int8 1000,000,000 100 100.0 260.0 74.0
DEEP100M float 100,000,000 96 38.4 26.0 7.4
SIFT100M uint8 100,000,000 128 12.8 26.0 7.4
SPACEV100M int8 100,000,000 100 10.0 26.0 7.4
We present experiments on three popular real-world datasets that originate from a diverse set of applications. We summarize the characteristics of these datasets along with the sizes of the generated graphs and compressed vectors in Table II. DEEP1B [46] is a collection of one billion image embeddings compressed to 96 dimensions. SIFT1B [29] represents the 128-dimensional SIFT (Scale-Invariant Feature Transform) descriptors of one billion images. The SPACEV1B dataset [13] encodes web documents and web queries sourced from Bing using the Microsoft SpaceV Superior Model. This dataset contains more than one billion points, and to be consistent with other billion-scale datasets in our test suite, we pick the first billion points in the dataset. The 100M versions of the three datasets are generated by taking the first hundred million points from their respective 1B datasets. We use a query batch size of 10,000 and run all the queries in a single batch, as in previous approaches [49, 21].
@@ VI-C Algorithm Parameters: BANG
Graph Construction. We run BANG search algorithm on Vamana graph, built using DiskANN [25]. We specify $R=64$ (maximum degree bound), $L=200$ (size of worklist used during build) and $\sigma$=1.2 (pruning parameter), following the recommendations of the original source. The sixth column in Table II shows the size of the resulting graph for each dataset.
Compression. We specify the parameters for PQ compression such that the generated compressed vectors fit entirely in the GPU global memory. For our setup, we empirically determined the largest value of $m$ to be 74. We use $m=74$ across all datasets for all our experiments, unless otherwise stated. Note that as a consequence, the compression ratio could vary (although $m$ remains the same) for different datasets depending upon their dimensionality. For example, for the 128-dimension SIFT1B dataset, the compression ratio is $74/128=0.57$, while for the 384-dimension DEEP1B dataset, the compression ratio is $74/384=0.19$. The last column in Table II shows the total size of PQ compressed vectors for each dataset.
Search. The search starts from the medoid, a node pre-determined during graph construction. We use the standard $k\mbox{-}recall@k$ [25] recall metric.We measure the throughput using the standard definition of Queries Per Second or QPS. The search parameter $t$ in Algorithm 2 represents the size of the worklist $\mathcal{L}$. The recall increases with $t$. The lower bound for $t$ is $k$, and the maximum value of $t$ is empirically set to 152 for our experiments. The Bloom filter (Section IV-C) is created to hold 399,887 entries, which we determined empirically. We tune the Bloom filter size to lower values in order to generate lower recall values (below the recall values achieved by using $t=k$). To increase statistical significance, we report the throughput averaged over five independent runs.
@@ VI-D Baselines
We compare BANG with the state-of-the-art open-source techniques that support billion-scale ANNS on a single GPU. Techniques like SONG [49] were found to work well on the million-scale datasets but could not be scaled up to the billion-scales. Therefore, we did not consider them.
GGNN [21]. GGNN performs GPU-based ANNS on a hierarchical $k$-NN graph. It works by dividing a large dataset into small shards that can be processed on multiple GPUs in parallel. GGNN provides configurable graph construction parameters viz.:number of out-going edges ($k$), number of disjoint points selected for subgraphs ($s$), number of levels in the hierarchical graph ($l$), slack factor ($\tau_{b}$) and number of refinement iterations ($r$). In the original work, eight GPUs were used for billion-scale datasets, but we configured GGNN to run with a single GPU. We use the recommended build parameters from the original source code repository [22] for most datasets; for others, we follow authors’ private email suggestions for optimal performance on our setup. We configure 1B dataset as {$k$:20, $s$:32, $l$:4, $\tau_{b}$:0.6, $r$:2, 16 shards} and 100M dataset as {$k$:24, $s$:32, $l$:4, $\tau_{b}$:0.5, $r$:2, 2 shards}.
(a) SIFT1B
(b) DEEP1B
(c) SPACEV1B
Figure 5: QPS-recall curves on billion-size datasets.
(a) SIFT100M
(b) DEEP100M
(c) SPACEV100M
Figure 6: QPS-recall curves on 100M datasets.
FAISS [27]. FAISS supports billion-scale ANN search on GPU. The underlying search structure of FAISS is the inverted-index (IVF) [34]: points are typically assigned to buckets via single-level clustering or two-level clustering. FAISS provides parameters for users to specify the type of quantization to be used on the base dataset. It supports two compression techniques, Scalar Quantization (SQ) and Product Quantization (PQ), and we perform experiments using each technique separately. We followed the guidelines mentioned in the FAISS Wiki [26] to experiment with different values for the IVF index type, and PQ/SQ transform to achieve optimal recall and throughput. For PQ, we configured the index string as: OPQ32_128, IVF262144, PQ32. The first field indicates the vector transformation technique (pre-processing) that aids in the subsequent steps of index construction and compression stages. The second field indicates the type of IVF index used. The third field indicates the encoding scheme (quantization) used. For SQ, we configured the index string as: IVF65536, SQ8 as indicated in Billion-Scale Approximate Nearest Neighbor Search Challenge [39]. Note that SQ indexes are less space-efficient than PQ but provide better recalls. All experiments were performed on a single GPU.

### v4

@@ IV-E Prefetching Candidate Nodes
As described in Algorithm 2, candidate node $u^{*}_{i}$ of each query $q_{i}$ will be communicated to the CPU (Line 2) only after the GPU has updated the worklist $\mathcal{L}_{i}$. After that, while the CPU fetches the neighbours (Line 2) and communicates them to the GPU (Line 2), the GPU remains idle, waiting to receive this information. To avoid this delay in starting the current iteration, GPU must be supplied with the neighbour IDs as soon as needed. To address this, instead of sending the candidate nodes at the end of the previous iteration to CPU, we perform an optimisation to eagerly identify the next candidate nodes immediately after the neighbour distance calculation (i.e., before even sorting the neighbour list by distance and updating worklist $\mathcal{L}_{i}$).
Our optimisation calculates the candidate node $u^{*}_{i}$ eagerly by selecting the nearest neighbour in the new neighbour list $N^{\prime}_{i}$ and the first unvisited node in worklist $\mathcal{L}_{i}$ (sorted by distance to query), then choosing the best between the two. The eager selection occurs just before Line 2 in Algorithm 2. Upon dispatching the eagerly selected candidates to the CPU, the GPU continues executing the remaining tasks of the iteration (sorting the new neighbour list and updating the worklist). Concurrently, the CPU fetches neighbour IDs. That is, Line 2 executes on CPU concurrently with tasks at Line 2 and 2 on GPU. By the completion of the GPU’s remaining tasks, the CPU has already communicated the neighbour IDs. This strategy eliminates GPU idle time, resulting in an increase in throughput and experimentally found to be $8-12\%$.
@@ IV-F Parallel Merge Sort
Work-Span Analysis. The work of merging two lists of size $t$ using the parallel merge routine is $O(t\cdot log(t))$ (Section IV-G). The total work $T(n)$ for an array of $n$ numbers follows the recurrence: $T(n)=2\cdot T(n/2)+n\cdot log(n)$, which gives $T(n)=O(n\cdot log^{2}(n))$. Thus, the work of the algorithm is $O(R\cdot log^{2}(R)\cdot\rho)$. Owing to the parallelisation scheme, the span of the algorithm is $O(log^{2}(R))$.
@@ IV-G Parallel Merge
The parallel merge routine that we detail in this section is employed in the parallel merge sort, as explained in the preceding section. Additionally, it is utilised to merge the sorted neighbour list $\mathcal{N}^{\prime}_{i}$ with the worklist $\mathcal{L}_{i}$ at Line 2 in Algorithm 2.
Figure 3 shows our parallel list merge algorithm on two lists of size four each. Consider two sorted lists $L_{in1}$ and $L_{in2}$ having $m$ and $n$ items respectively. Suppose that an item $e$ in $L_{in1}$ is at position $p_{1}<m$, and if inserted in $L_{in2}$ while maintaining sorted order, it will be placed at position $p_{2}<n$. Then, the position $pos_{Lout}$ of the element in the merged list $L_{out}$ is determined by $pos_{Lout}\leftarrow p_{1}+p_{2}$. We assign one thread to each element. From the thread ID assigned to the item, we can compute the position of the item in its original list. We use binary search to determine where the item should appear in the other list. For example, in Figure 3, the item 28 in $L_{in1}$ is at index 3, determined by the thread ID. In $L_{in2}$, its index is 4, found through binary search. Therefore, the index of 28 in the merged list is 7 (i.e., 3 + 4). Each thread, assigned to an item, concurrently performs a binary search for that item in the other list. We keep the lists in the GPU’s shared memory to minimise search latency. Thus, computing the position of each element in the merged list takes $O(\log(m)+\log(n))$ time. Lastly, elements are scattered to their new positions in the merged list; in the example, 28 is placed at index 7 in the merged list $L_{out}$. This step is also performed in parallel, as each thread writes its item to its unique position in the merged list.
Work-Span Analysis. The work of the algorithm to merge two lists each of size $\ell$ is $O(\ell\cdot log(\ell))$. Owing to the parallelisation scheme, the span of the algorithm is $O(log(\ell))$.
@@ IV-H Re-ranking
We implement the re-ranking step as a separate kernel with optimised thread-block sizes. Re-ranking involves computing the exact (i.e. Euclidean) distance for each query vector with its respective candidate nodes, followed by sorting these candidates by their distance to the query vector and reporting the top-$k$ candidates as nearest neighbours. For a query, its distance from each candidate node is computed in parallel. To sort the candidates according to their exact distance, we use the parallel merge procedure described in Section IV-F. Thanks to the asynchronous data transfers (Section IV-B), the candidates’ base dataset vectors are already available on GPU when this kernel begins. From our experiments, we observed that re-ranking improved the recall by 10-15% for the datasets under consideration.
@@
$O((d\cdot|\mathcal{C}|+|\mathcal{C}|\cdot log^{2}(|\mathcal{C}|))\cdot\rho)$, where $|\mathcal{C}|$ is the maximum number of candidate nodes for a query and $d$ is vector dimension. Owing to the parallelisation scheme, the span of re-ranking is $O(d+log^{2}(|\mathcal{C}|))$.
@@ V Implementation Variants
We refer to the implementation described so far in Sections III and IV as BANG. In addition, we provide versions of the implementation specifically optimised for medium-sized datasets (having up to 100-million points) that can fit in the GPU memory along with their respective indexes. We discuss these versions below.
@@ V-A In-memory Version
Since the CPU-GPU interconnect stays constantly busy when we store the graph on the CPU, as in BANG, careful task management on the GPU is needed to avoid idle time due to CPU-GPU interdependency. However, following the same approach is unnecessary for smaller graphs that can fit on a GPU. Keeping the entire graph on the GPU, rather than the CPU, eliminates the memory-intensive process of the CPU fetching neighbours from its memory for GPU-supplied candidates, resulting in enhanced throughput. We refer to this version of the implementation as BANG In-memory. The schematic block diagram of this version is shown in Figure 4. There are two main differences from Figure 1: (1) in step 1, we transfer the entire graph from CPU to GPU before the search begins (2) steps 2 and 3 fetch neighbours on the GPU locally through accesses to global memory instead of relying on the CPU.
@@ V-B Exact-distance Version
This variant further optimises the BANG In-memory version. Instead of using PQ distances within the greedy search loop and subsequently compensating the inaccuracies (resulting due to the use of PQ distances) by the re-ranking step, we directly use exact (Euclidean) distance calculations instead of PQ distances. We refer to this version as BANG Exact-distance. The main differences from Figure 4 are: (1) absence of PQ Distance Table computation, (2) exact distance computation (instead of PQ distance summation), (3) absence of re-ranking step.
@@ VI-A Machine Configuration
Hardware. We conduct our experiments on a machine with 2 x Intel Xeon Gold 6326 CPU (with 32 cores), clock-speed 2.90GHz, and 660 GiB DDR4 memory. (For the Text-to-Image-1B dataset, we used 1133 GiB due to its large size.) It houses a 1.41 GHz NVIDIA Ampere A100 GPU with 80 GB global memory and a peak bandwidth of 2039 GB/s. The GPU is connected to the host via a PCIe Gen 4.0 bus having a peak transfer bandwidth of 32 GB/s. We use a single GPU for all the experiments.
Software. The machine runs Ubuntu 22.04.01 (64-bit). We use g++ version 11.3 with -O3, -std=c++11 and -fopenmp flags to compile the C++ code. We use OpenMP for parallelising the C++ code. The GPU CUDA code is compiled using nvcc version 11.8 with the -O3 flag. The source code of BANG is publicly available at https://github.com/karthik86248/BANG-Billion-Scale-ANN.
@@ VI-B Datasets
TABLE II: Real-world datasets in our test-suite. (CV Size stands for Compressed Vectors size)
Number of Vector Graph CV
Dataset Type Data Points $d$ Size(GB) Size(GB) Size(GB)
SIFT-1B [33] uint8 1000,000,000 128 128.0 260.0 74.0
Deep-1B [62] float 1000,000,000 96 384.0 260.0 74.0
MSSPACEV-1B [14] int8 1000,000,000 100 100.0 260.0 74.0
Text-to-Image-1B [17] float 1000,000,000 200 800.0 260.0 74.0
SIFT-100M uint8 100,000,000 128 12.8 26.0 7.4
Deep-100M float 100,000,000 96 38.4 26.0 7.4
MSSPACEV-100M int8 100,000,000 100 10.0 26.0 7.4
Text-to-Image-100M float 100,000,000 200 80.0 26.0 7.4
We present experiments on four popular real-world datasets that originate from a diverse set of applications. We summarise the characteristics of these datasets along with the sizes of the graphs generated using Vamana and compressed vectors in Table II. The Deep dataset [62] is a collection of one billion image embeddings compressed to 96 dimensions. The SIFT dataset [33] dataset represents the 128-dimensional SIFT (Scale-Invariant Feature Transform) descriptors of one billion images. The Microsoft SPACEV (MSSPACEV) dataset [14] encodes web documents and web queries sourced from Bing using the Microsoft SpaceV Superior Model. This dataset contains more than one billion points, and to be consistent with other billion-scale datasets in our test suite, we pick the first billion points in the dataset. The Text-to-Image (T2I) dataset, derived from the Yandex [17] visual search engine, comprises one billion 200-dimensional image embeddings for indexing. It focuses on a cross-domain setting, where a user provides a textual query, and the search engine retrieves the most relevant images. This scenario introduces the possibility of differing distribution patterns between the dataset and the query vectors, also referred to as an Out-Of-Distribution (OOD) dataset, which makes it challenging to handle. Henceforth, for brevity, we refer to datasets containing one billion data points as 1B datasets and those with 100 million data points as 100M datasets. The 100M versions of the four datasets are generated by extracting the first hundred million points from the respective 1B datasets. We use a query batch size of 10,000 unless otherwise stated, with all queries executed concurrently in a single batch, following prior works [67, 24]. For datasets that provide more than 10,000 queries, we select a subset consisting of 10,000 queries.
@@ VI-C Algorithm Parameters: BANG
Graph Construction. We run BANG search algorithm on Vamana graph, built using DiskANN [29]. We specify $R=64$ (maximum degree bound), $L=200$ (size of worklist used during build) and $\sigma$=1.2 (pruning parameter), following the recommendations of the original source. The rightmost column in Table II shows the size of the resulting graph for each dataset.
Bloom filter. We use a Bloom filter per query, of size $\sim 400$ kB. Our Bloom filter uses an array of $z$ bools, where $z$ is determined using an estimate of the number of processed nodes for a query, a small tolerable false-positive rate, and the number of hash functions used. We use two FNV-1a hash functions [23], which are lightweight, non-cryptographic hash functions often used to implement Bloom filters.
Compression. During graph construction, we configure the parameters for PQ compression such that the generated compressed vectors fit entirely in the GPU global memory. For our setup, we empirically determined the largest value of $m$ to be 74. We use $m=74$ across all datasets for all our experiments unless otherwise stated. Note that as a consequence, the compression ratio could vary (although $m$ remains the same) for different datasets depending upon their dimensionality. For example, for the 128-dimension SIFT-1B dataset, the compression ratio is $74/128=0.57$, while for the 384-dimension Deep-1B dataset, the compression ratio is $74/384=0.19$. The rightmost column in Table II shows the total size of PQ compressed vectors for each dataset.
Search. The search starts from the medoid, a node pre-determined during graph construction. We use the widely used $k\mbox{-}recall@k$ [29] recall metric. The goal of ANNS is to efficiently retrieve, for a query point $q$, a set $\tilde{S}$ of $k$ candidate points from the dataset in order to maximise $k$-$recall@k\coloneq\frac{|S\cap\tilde{S}|}{k}$, where $S$ is the ground truth of the $k$ nearest neighbours of $q$ in the dataset. We measure the throughput using Queries Per Second or QPS which is (number of queries processed / search time), where search time is the time taken to process all queries. The search parameter $t$ in Algorithm 2 represents the size of the worklist $\mathcal{L}$. The recall increases with $t$. The lower bound for $t$ is $k$, and the maximum value of $t$ is empirically set to 152 for our experiments. The Bloom filter (Section IV-C) is created to hold 399,887 entries per query, which we determined empirically. We tune the Bloom filter size to lower values in order to generate lower recall values (below the recall values achieved by using $t=k$). To increase statistical significance, we report the average throughput over five independent runs.
@@ VI-D Baselines
We primarily evaluate BANG against the state-of-the-art open-source ANNS methods GGNN [24] and FAISS [31] that support billion-scale data on a single GPU. While our focus is on billion-scale datasets, we include comparisons on 100M datasets for completeness, including comparison with CAGRA [46], the leading GPU-based ANNS method for 100M points, despite its lack of support for billion-scale data.
GGNN [24]. GGNN provides configurable graph construction parameters viz. number of out-going edges ($k$), number of disjoint points selected for subgraphs ($s$), number of levels in the hierarchical graph ($l$), slack factor ($\tau_{b}$) and number of refinement iterations ($r$). In the original work, eight GPUs were used for billion-scale datasets, but we configured GGNN to run with a single GPU. We use the recommended build parameters from the original source code repository [25] for most datasets; for others, we follow authors’ private email suggestions for optimal performance on our setup. We configure 1B dataset as {$k$:20, $s$:32, $l$:4, $\tau_{b}$:0.6, $r$:2, 16 shards} and 100M dataset as {$k$:24, $s$:32, $l$:4, $\tau_{b}$:0.5, $r$:2, 2 shards}.
(a) SIFT-1B
(b) Deep-1B
(c) MSSPACEV-1B
(d) Text-to-Image-1B
Figure 5: Throughput (y-axis) v/s 10-recall@10 for 1B datasets. y-axis is in log-scale.
(a) SIFT-1B
(b) Deep-1B
(c) MSSPACEV-1B
(d) Text-to-Image-1B
Figure 6: Throughput (y-axis) v/s 100-recall@100 for 1B datasets. y-axis is in log-scale.
(a) SIFT-100M
(b) Deep-100M
(c) MSSPACEV-100M
(d) Text-to-Image-100M
Figure 7: Throughput (y-axis) v/s 10-recall@10 for 100M datasets. y-axis is in log-scale.
FAISS [31]. FAISS provides users with configuration options for quantizing the base dataset using either Scalar Quantization (SQ) or Product Quantization (PQ); following the FAISS Wiki guidelines [30], we experiment with different IVF index types and PQ/SQ transforms to optimise recall and throughput. For PQ, we configured the index string as: OPQ32_128, IVF262144, PQ32. The first field specifies the vector transformation used for preprocessing, the second denotes the IVF index type, and the third indicates the quantization scheme applied. For SQ, we configured the index string as: IVF65536, SQ8 as indicated in Billion-Scale Approximate Nearest Neighbor Search Challenge [52]. Note that SQ indexes are less space-efficient than PQ but provide better recalls. We performed all experiments on a single GPU.
CAGRA [46]. CAGRA is a state-of-the-art ANNS algorithm in the Nvidia cuVS [43], a high-performance, GPU-accelerated library for efficient vector search. We deployed CAGRA using the RAFT’s ANN benchmark collection. We configured CAGRA with the NN-Descent [19] algorithm and set the kNN graph degree to 32. For optimal search results, we vary the internal priority queue size (itopk parameter) from 32 to 512.